OpenStudy (anonymous):

use integration by parts: integral of 2^(sqrtx)

OpenStudy (anonymous):

\[\int\limits_{4}^{9} 2^{\sqrt{x}} dx\] \[u= 2^{\sqrt{x}}\] \[du= 2^{\sqrt{x}} (\ln2)(\frac{ 1 }{ 2\sqrt{x}})\] \[dv= 1\] \[v= x\] \[= x(2^{\sqrt{x}}) - \frac{ \ln2 }{ 2 } \int\limits 2^{\sqrt{x}}(\sqrt{x})\]

OpenStudy (anonymous):

is the substitution given ?

OpenStudy (anonymous):

no, but i have to use Integration by parts..

OpenStudy (anonymous):

\[u= 2^{\sqrt{x}}\] \[du= 2^{x}(\ln2)(\frac{ 1 }{ 2x }) dx\]\[dv= \sqrt{x}\] \[v=\frac{ 2 x ^{\frac{ 3 }{ 3 }}}{ 3 }\] is this substitution right?

OpenStudy (anonymous):

\[v ={\frac{2x^{\frac{ 3 }{ 2 }}}{3}} \]

OpenStudy (alekos):

This is a tricky one. I know the answer but I don't know how it's done. give me 24 hours

OpenStudy (anonymous):

i'm trying to solve it but i'm still writing it in LaTex to post here

OpenStudy (anonymous):

d/dx a^x = a ln a

OpenStudy (anonymous):

a^x ln a*

OpenStudy (loser66):

Manipulate a little bit let u = 2^(sqrt x) --> \(log_2 u = \sqrt x\)--> x = (\(\log_2 u)^2 \)

OpenStudy (loser66):

try this way you can get the answer :)

OpenStudy (anonymous):

sorry i don't get it. but thanks :)

OpenStudy (solomonzelman):

§ 2\(\normalsize\color{slate}{ ^{\sqrt{x}} }\) dx u= \(\normalsize\color{slate}{ \sqrt{x} }\) du=1 / 2\(\normalsize\color{slate}{ \sqrt{x} }\) 2§ 2\(\normalsize\color{slate}{ ^{u} }\) u du f=u ds=2\(\normalsize\color{slate}{ ^{u} }\) du df=du s= 2\(\normalsize\color{slate}{ ^{u} }\) / log(2) 2§ 2\(\normalsize\color{slate}{ ^{u+1} }\) u 2§ 2\(\normalsize\color{slate}{ ^{u+1} }\) --------- - ---------- +C log(2) log(2)

OpenStudy (solomonzelman):

Damn these question marks... os always has issues -:(

OpenStudy (solomonzelman):

the integral signs shouldn't be there in a last line, my bad

OpenStudy (solomonzelman):

this is what it should be, without 2§ in the begging of each fraction.